OpenStudy (anonymous):

2cos^2x/2tanx = tanx + sinx

OpenStudy (anonymous):

how do you prove this

OpenStudy (anonymous):

So you're asking how to prove: \[\frac{ 2\cos^2(x) }{ 2\tan(x) } = \tan(x)+\sin(x)\] ?